the area of the playground is 96 square yards. complete the square to find the value of x. the length of the playground is x+15 and the width is x+5.
Answer by nerdybill(7384)

the area of the playground is 96 square yards. complete the square to find the value of x. the length of the playground is x+15 and the width is x+5.
.
96 = (x+15)(x+5)
FOIL right side:
96 = x^2+20x+75
0 = x^2+20x-21
21 = (x^2+20x)
21+100 = (x^2+20x+100)
121 = (x+10)^2
.
Now find square root both sides:
11 = x+10
1 = x
.
length:
x+15 = 1+15 = 16 yards
width:
x+5 = 1+5 = 6 yards